A student has a 2.28 Liter bottle that contains a mixture of O2, N2, and CO2, with a total pressure of 5.70 bar at 298K . She knows that the mixture contains 0.221 mol N2 and that the partial pressure of CO2 is 0.290 bar . Calculate the partial pressure of O2.
Answer #1

Given:

P = 5.7 bar

= (5.7/1.01325) atm

= 5.6255 atm

V = 2.28 L

T = 298.0 K

find number of moles using:

P * V = n*R*T

5.6255 atm * 2.28 L = n * 0.08206 atm.L/mol.K * 298 K

n = 0.5245 mol

This is total mol of gas.

Then,

p(N2) = mole fraction of N2 * total pressure

= (0.221/0.5245) * 5.70 bar

= 2.40 bar

Now use:

p(O2) = Total pressure - p(N2) + p(CO2)

= 5.70 bar - 2.40 bar - 0.290 bar

= 3.01 bar

Answer: 3.01 bar